OpenStudy (anonymous):

Circle S has an equation of (x + 16)2 + (y – 9)2 = 4. What is the center and radius of circle S? Center: (16, -9); Radius: 4 Center: (16, -9); Radius: 2 Center: (-16, 9); Radius: 4 Center: (-16, 9); Radius: 2

OpenStudy (anonymous):

Equation of circle is (x+16)^2+(y-9)^2=4 [x-(-16)]^2+(y-9)^2=2^2 centre is (-16,9) and radius is 2

OpenStudy (anonymous):

thanks